OpenStudy (anonymous):

simplify (-6t)^3, writing the resulting expression with only positive exponents

OpenStudy (anonymous):

1/(6t)^3 i believe

OpenStudy (anonymous):

-6^3t^3=-36(6)t^3=-216t^3